# The Effect of Omitted Variables on

### v3

@@ 8 Conclusion
In this paper we distinguished between two kinds of breakdown points: sign change and explain away. Using theoretical results that apply regardless of the data generating process (section 3), four detailed empirical applications (section 6 and appendix LABEL:sec:extraEmpirical), and two meta-analyses covering 58 papers (section 7), we showed that the distinction between these two types of breakdown points matters in most empirical settings. In particular, explain away breakdown points often substantially overstate the robustness of results to omitted variables that can flip their sign. Given these results, we recommend that researchers plot the estimated identified set for the coefficient of interest under a variety of different assumptions about omitted variables; for example, see figure 2. We also recommend reporting sign change breakdown points as robustness summary statistics, such as in table 6. Both of these recommendations are easily implemented by using our companion Stata package regsensitivity.

### v4

@@ 8 Conclusion
In this paper we distinguished between two kinds of breakdown points: sign change and explain away. Using theoretical results that apply regardless of the data generating process (Section 3), four detailed empirical applications (Section 6 and Appendix LABEL:sec:extraEmpirical_applications), and two meta-analyses covering 58 papers (Section 7), we showed that the distinction between these two types of breakdown points matters in most empirical settings. In particular, explain away breakdown points often substantially overstate the robustness of results to omitted variables that can flip their sign. Given these results, we recommend that researchers plot the estimated identified set for the coefficient of interest under a variety of different assumptions about omitted variables; for example, see Figure 2. We also recommend reporting sign change breakdown points as robustness summary statistics, such as in Table 6. Both of these recommendations are easily implemented by using our companion Stata package regsensitivity.